OpenStudy (anonymous):

Part 1: Solve the inequality 12 + x <= 2x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

subtract x from both sides, then add 2 to both sides, then divide by 2 on both sides, then tell me what you got

OpenStudy (anonymous):

4<=x @zzr0ck3r

OpenStudy (anonymous):

Then for the second one i got x>=2 @zzr0ck3r

OpenStudy (zzr0ck3r):

12 + x <= 2x - 2 subtract x 12<=x-2 add 2 14<=x

OpenStudy (anonymous):

oh i see so whats next

OpenStudy (zzr0ck3r):

thats it, solve means get x alone on one side.

OpenStudy (anonymous):

So now whats the final answer

OpenStudy (zzr0ck3r):

14<=x

OpenStudy (anonymous):

so do you combine the first and second 1 @zzr0ck3r

OpenStudy (zzr0ck3r):

there is only 1.

OpenStudy (zzr0ck3r):

show me what you did to get your answer

OpenStudy (anonymous):

So for part 1 its 14<=x

OpenStudy (anonymous):

@texaschic101 can you help me out

OpenStudy (zzr0ck3r):

graph it its the x axis with all x>= 14

OpenStudy (anonymous):

So thats the final answer

OpenStudy (texaschic101):

12 + x <= 2x - 2 12 + 2 <= 2x - x 14 <= x Are you using a number line to graph ? If so, closed circle on 14, shaded to the left

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

So whats the final answer for part 1

OpenStudy (texaschic101):

oops...closed circle on 14, shaded to the right. <--the graph The answer to part 1 is : 14 <= x

OpenStudy (texaschic101):

if it is greater then you shade to the right, if it is less then, you shade to the left. We used a closed circle because an equal sign was involved.

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

I have told you the answer to part 1 twice now:)